# Joint Discriminative-Generative Modeling via Dual Adversarial Training

### v1

@@ A.1 Extended discussion on related work
A.2 Intuitive connection between $R_{1}$ regularization and adversarial training
We provide mathematical intuition for why adversarial training can serve as implicit regularization in place of explicit $R_{1}$ gradient penalties. For classification tasks, consider a vector-valued function $f:\mathbb{R}^{d}\to\mathbb{R}^{K}$ producing logits for $K$ classes.
@@ A.2 Intuitive connection between $R_{1}$ regularization and adversarial training
$\mathcal{L}_{\text{AT}}=\mathbb{E}_{(x,y)\sim p_{\text{data}}}\left[\max_{\|\delta\|_{2}\leq\epsilon}L(f(x+\delta),y)\right]$ (15)
To understand how adversarial training provides implicit gradient regularization, we analyze the first-order behavior of this cross-entropy adversarial objective.
First-order expansion of the adversarial objective.
Let $z=f(x)\in\mathbb{R}^{K}$, $p=\text{softmax}(z)$, and $J_{f}(x)\in\mathbb{R}^{K\times d}$ be the input-Jacobian of the logits. For cross-entropy loss $L(z,y)=-\log p_{y}$, a first-order Taylor expansion gives:
$L(f(x+\delta),y)$ $=L(f(x),y)+\nabla_{x}L(f(x),y)^{T}\delta+O(\|\delta\|_{2}^{2})$ (16)
$\max_{\|\delta\|_{2}\leq\epsilon}L(f(x+\delta),y)$ $\approx L(f(x),y)+\epsilon\|\nabla_{x}L(f(x),y)\|_{2}$ (17)
This approximation is valid for sufficiently small $\epsilon$ relative to the local curvature of $L$, such that higher-order terms remain negligible over the $L_{2}$ constraint ball $\{\delta:\|\delta\|_{2}\leq\epsilon\}$.
This shows that adversarial training implicitly adds a penalty term proportional to $\|\nabla_{x}L\|_{2}$ (the first power of the gradient norm). To understand what this gradient represents, we apply the chain rule:
$\nabla_{x}L=J_{f}(x)^{T}\nabla_{z}L=J_{f}(x)^{T}(p-e_{y})$ (18)
where $e_{y}$ is the one-hot label vector. Let $g_{k}(x):=\nabla_{x}f_{k}(x)\in\mathbb{R}^{d}$ denote the per-class input gradients (the rows of $J_{f}$). Then we can write:
$\nabla_{x}L=\sum_{k=1}^{K}(p_{k}-\delta_{ky})g_{k}=-(1-p_{y})g_{y}+\sum_{k\neq y}p_{k}g_{k}$ (19)
This decomposition reveals that the cross-entropy gradient is a weighted combination of per-class gradients, where the true class gradient $g_{y}$ appears with negative weight $(1-p_{y})$ and competitor gradients $g_{k}$ appear with positive weights $p_{k}$.
Expansion into $R_{1}$-style components.
To understand how this relates to standard $R_{1}$ regularization, we can expand the squared gradient norm by substituting the final expression for $\nabla_{x}L$. Although the actual adversarial penalty is proportional to $\|\nabla_{x}L\|_{2}$, examining $\|\nabla_{x}L\|_{2}^{2}$ provides useful analytical insight:
$\|\nabla_{x}L\|_{2}^{2}=\left\|-(1-p_{y})g_{y}+\sum_{k\neq y}p_{k}g_{k}\right\|_{2}^{2}$ (20)
Expanding this expression:
$\|\nabla_{x}L\|_{2}^{2}$ $=\underbrace{(1-p_{y})^{2}\|g_{y}\|_{2}^{2}}_{\text{down-weighted true-class }R_{1}}+\underbrace{\sum_{k\neq y}p_{k}^{2}\|g_{k}\|_{2}^{2}}_{\text{competitor }R_{1}\text{ terms}}$
$\quad\underbrace{-2(1-p_{y})\sum_{k\neq y}p_{k}\langle g_{y},g_{k}\rangle}_{\text{true vs competitor alignment}}+\underbrace{2\sum_{\begin{subarray}{c}i<j\\ i,j\neq y\end{subarray}}p_{i}p_{j}\langle g_{i},g_{j}\rangle}_{\text{competitor-competitor alignment}}$ (21)
This expansion decomposes the adversarial penalty into interpretable components:
  1. 1.
    True-class $R_{1}$ regularization: $(1-p_{y})^{2}\|g_{y}\|_{2}^{2}$, which is the standard $R_{1}$ penalty on the true class, down-weighted by confidence
  2. 2.
    Competitor $R_{1}$ terms: $\sum_{k\neq y}p_{k}^{2}\|g_{k}\|_{2}^{2}$, providing $R_{1}$-style regularization on competing classes weighted by their predicted probabilities
  3. 3.
    Gradient alignment terms: Cross-class inner products $\langle g_{i},g_{j}\rangle$ that discourage competitor–competitor alignment (favoring orthogonality), while encouraging the true-class gradient to align with the competitor average
The above decomposition shows that $\mathbb{E}_{(x,y)}[\|\nabla_{x}L\|_{2}^{2}]$ includes the core $R_{1}$ regularization term $(1-p_{y})^{2}\|g_{y}\|_{2}^{2}$ alongside additional smoothness constraints, suggesting that adversarial training may provide comparable regularization to explicit $R_{1}$ penalties in contexts where gradient control is important. However, there are important distinctions: adversarial training down-weights high-confidence points through the $(1-p_{y})^{2}$ factor, has competitor $R_{1}$ terms, and introduces cross terms that encourage specific gradient alignment patterns absent in standard $R_{1}$. Additionally, the analysis relies on the first-order approximation being valid, which may not hold for the $\epsilon$ values commonly used in practice, and the actual adversarial penalty is proportional to $\|\nabla_{x}L\|_{2}$ rather than its square. Despite these caveats, this mathematical framework provides insights for why adversarial training’s comprehensive implicit regularization may offer sufficient smoothness constraints for stable energy-based model training, without requiring explicit gradient penalties.
A.3 DAT training algorithm
@@
Although both approaches yield unbiased estimates, we find ancestral sampling to be practically superior for training stability, possibly because it leverages the classifier’s existing strong class representations to provide better mode coverage and mixing properties, while direct sampling from the marginal distribution often diverges. We also find ancestral sampling (conditional generation) yields substantially better FID than directly sampling from marginal distribution (see Table˜11).
@@
A.4 Model training
A.4.1 Training setup
@@
For Stage 1, we utilize a pretrained CIFAR-10 model from RATIO [Augustin et al., 2020] and pretrained ImageNet models from Salman et al. [2020], while training our own CIFAR-100 model following the RATIO methodology with the hyperparameters specified in Table 3. We select the EMA model with the best robust test accuracy as the final Stage 1 model.
For Stage 2, we initialize from the Stage 1 model and continue training with batch normalization disabled by setting all BN modules to evaluation mode. During this stage, we optimize the complete objective function $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$ using fixed learning rates as specified in Table˜3. The discriminative component $\mathcal{L}_{\text{AT-CE}}(\theta)$ continues to use the same adversarial settings as Stage 1 (see Table˜4), while the generative component $\mathcal{L}_{\text{BCE}}(\theta)$ employs the parameters detailed in Table˜5.
@@
Table 3: Training hyperparameters for both stages.
CIFAR-10/100 ImageNet Architecture WideResNet-34-10 ResNet-50/WideResNet-50-4 Optimizer SGD with Nesterov (momentum=0.9) SGD with Nesterov (momentum=0.9) Weight decay $5\times 10^{-4}$ $1\times 10^{-4}$ (Stage 1), $5\times 10^{-4}$ (Stage 2) Batch size 128 512 EMA [Gowal et al., 2020] Yes Yes Learning rate (Stage 1) 0.1 (cosine schedule, 300 epochs) 0.1 (step decay at epochs 30, 60, 90) Learning rate (Stage 2) 0.001 (CIFAR-10), 0.01 (CIFAR-100) 0.001 Batch normalization (Stage 1) Enabled (training mode) Enabled (training mode) Batch normalization (Stage 2) Disabled (eval mode) Disabled (eval mode)
Table 4: Adversarial training parameters for $\mathcal{L}_{\text{AT-CE}}$ (identical across Stage 1 and Stage 2).
CIFAR-10/100 ImageNet
PGD steps 10 2
PGD step size 0.1 2.0
$L_{2}$ perturbation bound 0.5 3.0
@@
CIFAR-10/100 ImageNet
Max PGD steps ($T$) 45 15/30/70
PGD step size 0.1 2.0
$L_{2}$ perturbation bound None (unconstrained) None (unconstrained)
OOD data source 80M Tiny Images [Torralba et al., 2008] Open Images [Krasin et al., 2016]
A.4.2 Data augmentation details
As described in Section˜3.6, we implement separate data augmentation pipelines for the discriminative and generative components of our objective function. Table˜6 summarizes these dataset-specific augmentation strategies. Note that augmentation strategies for $\mathcal{L}_{\text{AT-CE}}$ are identical to those used by RATIO [Augustin et al., 2020] for CIFAR-10/CIFAR-100, and identical to Salman et al. [2020] for ImageNet. The effects of CIFAR-10 augmentations are illustrated in Figure˜2.
Figure˜3 illustrates CIFAR-10 training curves from Stage 2 joint training with various augmentation strategies applied to $\mathcal{L}_{\text{BCE}}$ (while consistently using AutoAugment with Cutout for $\mathcal{L}_{\text{AT-CE}}$). Interestingly, the choice of augmentation for the generative component influences discriminative performance as well, as evidenced by the decline in robust test accuracy when using no augmentation. The best FID performance is achieved by no augmentation and random cropping with padding, which minimally distort the underlying data distribution $p_{\text{data}}$. Overall we find random cropping with padding provides the optimal balance between discriminative and generative performances.

### v3

@@ A.1 Extended related work
A.2 Theoretical characterization
We provide a formal characterization of what distribution our joint objective (Equation˜9) learns. Our analysis builds on the theoretical framework from Yin et al. [2022] (AT-EBM) and extends it to the conditional modeling setting by deriving the optimal class logits under the joint objective.
Optimal solution for class logits $f_{\theta}^{*}(x)[y]$.
Following Yin et al. [2022], the generative component of our objective can be expressed as a maximin optimization problem. Let $D(x)=\sigma(-E_{\theta}(x))$ where $E_{\theta}(x)=-\log\sum_{y}\exp(f_{\theta}(x)[y])$ is the marginal energy function. For the theoretical analysis below, we assume (1) the PGD attack in Equation˜8 converges to the global minimum of $E_{\theta}(x)$, (2) the model has sufficient capacity, and (3) infinite training data from $p_{\text{data}}(x,y)$. Under these assumptions, minimizing $\mathcal{L}_{\text{BCE}}(\theta)$ implicitly solves:
$\max_{D}\min_{p_{T}}U(D,p_{T})=\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]+\mathbb{E}_{x\sim p_{T}}[\log(1-D(x))]$ (14)
where $p_{T}$ represents the distribution of samples after PGD attack initialized from the auxiliary out-of-distribution dataset $p_{\text{ood}}$.
Under these assumptions, the optimal solution to Equation˜14 is characterized by Proposition 1 of Yin et al. [2022], which shows that at optimum $U(D^{*},p_{T}^{*})=-\log(4)$ with:
  1. 1.
    $D^{*}(x)=\frac{1}{2}$ for all $x\in\mathrm{Supp}(p_{\text{data}}(x))$
  2. 2.
    $D^{*}(x)\leq\frac{1}{2}$ for all $x\notin\mathrm{Supp}(p_{\text{data}}(x))$
  3. 3.
    $p_{T}^{*}$ is supported on $\{x:D(x)=\frac{1}{2}\}$
where $\mathrm{Supp}(p_{\text{data}}(x))$ denotes the support of the marginal data distribution.
The above result characterizes only the marginal energy $E_{\theta}(x)$, leaving the individual class logits $f_{\theta}(x)[y]$ underdetermined. We now derive their optimal values by incorporating the discriminative objective $\mathcal{L}_{\text{AT-CE}}$.
Proposition A.1 (Optimal class logits).
Under the assumptions stated above, at the optimal solution to the joint objective $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$, the class logits satisfy on the support:
$f_{\theta}^{*}(x)[y]=\log p_{\text{data}}(y|x)\quad\text{for all }x\in\mathrm{Supp}(p_{\text{data}}(x))$ (15)
Proof.
From the AT-EBM result above, on the support we have $D^{*}(x)=\sigma(-E_{\theta}^{*}(x))=\frac{1}{2}$, which implies:
$E_{\theta}^{*}(x)=0\implies-\log\sum_{y}\exp(f_{\theta}^{*}(x)[y])=0\implies\sum_{y}\exp(f_{\theta}^{*}(x)[y])=1$ (16)
The conditional distribution is defined as:
$p_{\theta}(y|x)=\frac{\exp(f_{\theta}(x)[y])}{\sum_{y^{\prime}}\exp(f_{\theta}(x)[y^{\prime}])}$ (17)
Substituting the constraint from Equation˜16:
$p_{\theta}^{*}(y|x)=\frac{\exp(f_{\theta}^{*}(x)[y])}{1}=\exp(f_{\theta}^{*}(x)[y])$ (18)
The adversarial cross-entropy objective $\mathcal{L}_{\text{AT-CE}}$ minimizes $\log p_{\theta}(y|x)$ over worst-case perturbations of $(x,y)\sim p_{\text{data}}(x,y)$. Under our capacity assumption, at optimality on the support this reduces to standard cross-entropy behavior, yielding $p_{\theta}^{*}(y|x)=p_{\text{data}}(y|x)$. Since the optimal predictive distribution naturally satisfies the normalization constraint $\sum_{y}p_{\text{data}}(y|x)=1$, the solution to the discriminative objective is fully compatible with the marginal energy constraint $\sum_{y}\exp(f_{\theta}^{*}(x)[y])=1$ derived from the generative objective. Therefore:
$\exp(f_{\theta}^{*}(x)[y])=p_{\text{data}}(y|x)$ (19)
Taking logarithms gives Equation˜15. ∎
On-support behavior.
Proposition A.1 implies that on the support: (1) the joint energy equals the negative conditional log-probability $E_{\theta}^{*}(x,y)=-f_{\theta}^{*}(x)[y]=-\log p_{\text{data}}(y|x)$, (2) the marginal energy is constant $E_{\theta}^{*}(x)=0$, and (3) the unnormalized marginal density is constant, $p_{\theta}^{*}(x)\propto\exp(-E_{\theta}^{*}(x))=1$. This confirms that the model assigns equal density to all points on the support, rather than learning the true data density $p_{\text{data}}(x)$. For datasets with deterministic labels where $p_{\text{data}}(y|x)=\delta_{y,y_{\text{true}}(x)}$, this implies $f_{\theta}^{*}(x)[y]=0$ for the true class and $f_{\theta}^{*}(x)[y]=$-$\infty$ otherwise.
Off-support behavior.
For $x\notin\mathrm{Supp}(p_{\text{data}}(x))$, the optimal solution to Equation˜14 constrains $D^{*}(x)\leq\frac{1}{2}$, which implies $E_{\theta}^{*}(x)\geq 0$ and thus $\sum_{y}\exp(f_{\theta}^{*}(x)[y])\leq 1$. Since $\mathcal{L}_{\text{AT-CE}}$ is only computed on the data support and its adversarial perturbations, the individual class logits are underdetermined for points far from the support. From the constraint $\sum_{y}\exp(f_{\theta}^{*}(x)[y])\leq 1$, we have $f_{\theta}^{*}(x)[y]\leq 0$ for all classes $y$, and consequently $E_{\theta}^{*}(x,y)=-f_{\theta}^{*}(x)[y]\geq 0$.
Comparison.
For datasets with deterministic labels, this creates a hierarchical energy structure:
  • Valid pairs $(x\in\mathrm{Supp},y=y_{\text{true}})$: The joint energy is exactly $E_{\theta}^{*}(x,y)=0$.
  • On-support, incorrect labels $(x\in\mathrm{Supp},y\neq y_{\text{true}})$: The joint energy diverges to $E_{\theta}^{*}(x,y)=\infty$.
  • Off-support (OOD) $(x\notin\mathrm{Supp},\text{any }y)$: The joint energy is bounded below, $E_{\theta}^{*}(x,y)\geq 0$, but the exact value is underdetermined.
Thus, the learned energy function $E_{\theta}(x,y)$ enables robust classification (via $\arg\min_{y}E_{\theta}(x,y)$), generation (via minimizing $E_{\theta}(x,y)$ over $x$), and OOD detection (via thresholding $\min_{y}E_{\theta}(x,y)$).
Finite-step vs. convergent PGD.
The theoretical analysis above assumes the PGD attack converges to the global minimum of $E_{\theta}(x)$, which leads to the maximin formulation in Equation˜14. In practice, however, we use finite-step PGD (Equation˜8), which does not explicitly solve the inner minimization problem, creating a crucial gap between theory and practice. Our experiments demonstrate that our approach achieves diverse, high-quality generation, suggesting that rather than converging to energy minima, finite-step PGD behaves like a sampler in practice: starting from different initializations in the OOD dataset, it explores different regions of the energy landscape, leading to diverse generated samples.
Summary and implications.
Our formal analysis reveals that our joint objective learns a fundamentally different quantity than MLE-based methods. While MLE-based JEM theoretically learns the full density $p_{\theta}(x)$ with a valid partition function (though in practice short-run SGLD fails to achieve this), our approach explicitly learns the support of the data distribution with $E_{\theta}(x)=0$ on the support. The optimal class logits $f_{\theta}^{*}(x)[y]=\log p_{\text{data}}(y|x)$ (Proposition A.1) reveal how the joint objective uniquely determines the solution on the support.
This support-based characterization has important implications. The constant marginal energy $E_{\theta}^{*}(x)=0$ on the support means the model assigns equal unnormalized density to all points on the support, theoretically discarding frequency information about the data distribution. This represents a significant theoretical limitation: the model cannot distinguish between common and rare examples within the support, and thus cannot perform density estimation tasks that require modeling relative frequencies.
Despite this limitation, the support-based formulation provides clear advantages: superior training stability compared to SGLD-based methods, robust classification, and effective OOD detection—benefits that do not require density information. Additionally, our strong empirical generation results demonstrate that the finite-step PGD dynamics discussed above act as an effective sampler, capturing density variations despite the theoretical prediction of uniformity. Overall, the support-based approach provides a practical and stable framework for joint modeling, trading full density estimation for superior training stability and strong empirical performance.
A.3 Implicit $R_{1}$ regularization
Building on the operator norm equivalence established by Roth et al. [2020], we show that adversarial training implicitly bounds the $R_{1}$ gradient penalty, eliminating the need for explicit gradient regularization. Consider a classifier $f:\mathbb{R}^{d}\to\mathbb{R}^{K}$ producing logits for $K$ classes.
@@ A.3 Implicit $R_{1}$ regularization
$\mathcal{L}_{\text{AT}}=\mathbb{E}_{(x,y)\sim p_{\text{data}}}\left[\max_{\|\delta\|_{p}\leq\epsilon}L(f(x+\delta),y)\right]$ (21)
Adversarial training as operator norm regularization.
Roth et al. [2020] proved that $\ell_{p}$-norm constrained adversarial training with an $\ell_{q}$-norm loss on the logits is equivalent to data-dependent $(p,q)$-operator norm regularization:
$\mathcal{L}_{\text{AT}}\equiv\mathcal{L}_{\text{standard}}+\lambda(\epsilon)\cdot\|J_{f}(x)\|_{p,q}$ (22)
where $\lambda(\epsilon)$ is an implicit regularization coefficient that scales linearly with the attack budget $\epsilon$. Here $J_{f}(x)\in\mathbb{R}^{K\times d}$ denotes the input-Jacobian of the logits, and the $(p,q)$-operator norm is defined as:
$\|J_{f}(x)\|_{p,q}:=\max_{\|v\|_{p}=1}\|J_{f}(x)v\|_{q}$ (23)
This measures the maximal signal amplification when propagating a norm-bounded input perturbation through the linearized network. For $\ell_{2}$-constrained adversarial training, the $(2,2)$-operator norm reduces to the spectral norm $\sigma_{\max}(J_{f}(x))$, the largest singular value of the Jacobian:
$\mathcal{L}_{\text{AT}}\equiv\mathcal{L}_{\text{standard}}+\lambda(\epsilon)\cdot\sigma_{\max}(J_{f}(x))$ (24)
The optimal perturbation direction aligns with the dominant right singular vector of $J_{f}(x)$.
Connection to $R_{1}$ regularization.
Since adversarial training regularizes $\sigma_{\max}(J_{f}(x))$, we trace how this bound propagates to $\|\nabla_{x}f_{y}(x)\|_{2}$, which $R_{1}$ penalizes. By definition, the spectral norm satisfies $\sigma_{\max}(J_{f}(x))=\sigma_{\max}(J_{f}(x)^{T})=\max_{\|v\|_{2}=1}\|J_{f}(x)^{T}v\|_{2}$. Let $e_{y}\in\mathbb{R}^{K}$ be the standard basis vector for class $y$. The product $J_{f}(x)^{T}e_{y}$ extracts the $y$-th column of $J_{f}(x)^{T}$, which corresponds to the gradient $\nabla_{x}f_{y}(x)$. Since $\|e_{y}\|_{2}=1$, we have:
$\|\nabla_{x}f_{y}(x)\|_{2}=\|J_{f}(x)^{T}e_{y}\|_{2}\leq\max_{\|v\|_{2}=1}\|J_{f}(x)^{T}v\|_{2}=\sigma_{\max}(J_{f}(x))$ (25)
Thus, adversarial training’s regularization of $\sigma_{\max}(J_{f}(x))$ upper-bounds the $R_{1}$ gradient penalty $\|\nabla_{x}f_{y}(x)\|_{2}$.
Bound on all per-class gradients.
This analysis reveals that adversarial training provides a stronger regularization than $R_{1}$ alone. While $R_{1}$ penalizes only $\|\nabla_{x}f_{y}(x)\|_{2}$, adversarial training simultaneously bounds all per-class gradient norms:
$\|\nabla_{x}f_{k}(x)\|_{2}\leq\sigma_{\max}(J_{f}(x)),\quad\forall k\in\{1,\ldots,K\}$ (26)
This includes the probability-weighted mean gradient $\sum_{k}p_{k}\nabla_{x}f_{k}(x)$, since its norm is bounded by $\max_{k}\|\nabla_{x}f_{k}(x)\|_{2}$.
Effective regularization strength scales with attack budget.
An important consequence of this framework is that the effective regularization strength scales with the adversarial attack budget $\epsilon$. From the equivalence established above, the adversarial training objective implicitly regularizes $\sigma_{\max}(J_{f}(x))$ with strength $\lambda(\epsilon)$. For the network to maintain bounded total loss, the regularization term must remain bounded: $\lambda(\epsilon)\cdot\sigma_{\max}(J_{f}(x))\lesssim C$ for some constant $C$. Since $\lambda(\epsilon)$ scales linearly with $\epsilon$, this implies:
$\sigma_{\max}(J_{f}(x))\lesssim\frac{C}{\lambda(\epsilon)}$ (27)
Combining with our direct bound:
$\|\nabla_{x}f_{y}(x)\|_{2}\leq\sigma_{\max}(J_{f}(x))\lesssim\frac{C}{\lambda(\epsilon)}$ (28)
Thus, the $R_{1}$ gradient norm scales inversely with the regularization strength. Larger attack budgets $\epsilon$ induce stronger implicit regularization $\lambda(\epsilon)$, leading to lower $R_{1}$ gradients.
Connection to GAN regularization.
GAN training commonly employs spectral normalization [Miyato et al., 2018] on weight matrices and $R_{1}$ gradient penalties [Mescheder et al., 2018] for stability. The weight spectral norm provides a global upper bound on the Jacobian spectral norm: $\sigma(J_{f}(x))\leq\prod_{\ell}\sigma(W^{\ell})$ [Yoshida and Miyato, 2017]. However, Roth et al. [2020] show this bound can be loose as it ignores data-dependent activation patterns. AT’s equivalence to data-dependent spectral norm regularization is more targeted, directly controlling sensitivity at actual data points. Combined with our analysis showing this bounds $R_{1}$, AT provides a principled implicit alternative to both explicit GAN regularizers.
Empirical validation.
We validate this theoretical analysis by tracking $R_{1}$ gradient norms during Stage 2 joint training on ImageNet $\times256$ (ResNet-50, $T=15$). Figure 2 compares two settings: (1) adversarial training on the discriminative loss ($\mathcal{L}_{\text{AT-CE}}$), and (2) standard training on the discriminative loss. As predicted by the analysis above, the adversarial training curve remains bounded and stable throughout training, while standard training exhibits gradient explosion—with $R_{1}$ values increasing by orders of magnitude in later stages. While the exact equivalence from Roth et al. [2020] holds under technical conditions (small $\epsilon$ relative to ReLU cell size, infinite attack step size limit), our empirical observations are consistent with the theoretical predictions. This confirms that implicit regularization from adversarial training suffices to maintain bounded $R_{1}$ gradients, thereby enabling stable energy-based model training.
Figure 2: $R_{1}$ gradient norm during Stage 2 joint training on ImageNet $\times256$ (ResNet-50, $T=15$). Y-axis is in log scale. Adversarial training maintains bounded gradients while standard training exhibits gradient explosion.
A.4 Training algorithm
@@
Although both approaches yield unbiased estimates, we find ancestral sampling to be practically superior for training stability, possibly because it leverages the classifier’s existing strong class representations to provide better mode coverage and mixing properties, while direct sampling from the marginal distribution often diverges. We also find ancestral sampling (conditional generation) yields substantially better FID than directly sampling from marginal distribution (see Table˜9).
@@
Appendix B Training & Evaluation
B.1 Implementation details
B.1.1 Hyperparameters
@@
For Stage 1, we utilize a pretrained CIFAR-10 model from RATIO [Augustin et al., 2020] and pretrained ImageNet ResNet-50 and WRN-50-4 models from Salman et al. [2020], while training our own CIFAR-100 model following the RATIO methodology with the hyperparameters specified in Table 3. For ConvNeXt-Large experiments on ImageNet, we use the pretrained robust checkpoint from Singh et al. [2023]. We select the EMA model with the best robust test accuracy as the final Stage 1 model.
For Stage 2, we initialize from the Stage 1 model and continue training. For ResNet and WRN architectures, we disable batch normalization by setting all BN modules to evaluation mode. During this stage, we optimize the complete objective function $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$ using fixed learning rates as specified in Table˜3. For the discriminative component $\mathcal{L}_{\text{AT-CE}}(\theta)$, CIFAR and ImageNet ResNet/WRN models continue to use the same adversarial settings as stage 1, while ConvNeXt transitions from $L_{\infty}$ to $L_{2}$ perturbations with adjusted PGD parameters (see Table˜4). The generative component $\mathcal{L}_{\text{BCE}}(\theta)$ employs the parameters detailed in Table˜5.
@@
Table 3: Training hyperparameters for both stages. Epochs for stage 2 are estimated based on the number of in-distribution images seen by the discriminative component during training.
CIFAR-10/100 ImageNet ImageNet Architecture WRN-34-10 ResNet-50/WRN-50-4 ConvNeXt-L-CvSt BatchNorm Enabled (stage 1), Disabled (stage 2) Enabled (stage 1), Disabled (stage 2) N/A LayerNorm N/A N/A Enabled Optimizer SGD with Nesterov SGD with Nesterov AdamW Weight decay $5\times 10^{-4}$ $1\times 10^{-4}$ (stage 1), $5\times 10^{-4}$ (stage 2) 0.05 Batch size 128 512 756 (stage 1), 512 (stage 2) EMA Enabled Enabled Enabled LR (stage 1) 0.1 (cosine schedule) 0.1 (step decay at epochs 30, 60, 90) 0.001 (cosine decay with warm-up) LR (stage 2) 0.001 (CIFAR-10), 0.009 (CIFAR-100) 0.001 0.0003 Epochs (stage 1) 300 90 100 Epochs (stage 2) 26 (CIFAR-10), 30 (CIFAR-100) 0.78 (ResNet-50), 0.39 (WRN-50-4) 1.09
Table 4: Adversarial training parameters for $\mathcal{L}_{\text{AT-CE}}$ (identical across stages for CIFAR and ImageNet ResNet/WRN).
CIFAR-10/100 ImageNet (ResNet/WRN) ImageNet (ConvNeXt-L-CvSt) PGD steps 10 2 2 PGD step size 0.1 2.0 2/255 ($L_{\infty}$, stage 1), 2.0 ($L_{2}$, stage 2) Perturbation bound $L_{2}$, $\epsilon=0.5$ $L_{2}$, $\epsilon=3.0$ $\epsilon=4/255$ ($L_{\infty}$, stage 1), $\epsilon=3.0$ ($L_{2}$, stage 2)
@@
CIFAR-10/100 ImageNet (ResNet/WRN) ImageNet (ConvNeXt-L-CvSt) Max PGD steps ($T$) 40/45/50 15/30/65 110 PGD step size 0.1 2.0 3.0 $L_{2}$ perturbation bound None (unconstrained) None (unconstrained) None (unconstrained) OOD data source 80M Tiny Images [Torralba et al., 2008] Open Images [Krasin et al., 2016] Open Images [Krasin et al., 2016]
B.1.2 Data augmentation
As described in Section˜3.6, we implement separate data augmentation pipelines for the discriminative and generative components of our objective function. Table˜6 summarizes these augmentation strategies for stage 2 training. For CIFAR-10/100 and ImageNet ResNet/WRN, the pretrained models we initialize from were trained with the same augmentations as stage 2 $\mathcal{L}_{\text{AT-CE}}$ (from Augustin et al. [2020] for CIFAR and Salman et al. [2020] for ImageNet). For ImageNet ConvNeXt-L-CvSt, the pretrained model from Singh et al. [2023] was trained with heavy augmentations (RandAugment + MixUp + CutMix + Random Erasing), which differ from the simpler stage 2 $\mathcal{L}_{\text{AT-CE}}$ strategy shown in Table˜6. The effects of CIFAR-10 augmentations are illustrated in Figure˜3.
Figure˜4 illustrates CIFAR-10 training curves from stage 2 joint training with various augmentation strategies applied to $\mathcal{L}_{\text{BCE}}$ (while consistently using AutoAugment with Cutout for $\mathcal{L}_{\text{AT-CE}}$). Interestingly, the choice of augmentation for the generative component influences discriminative performance as well, as evidenced by the decline in robust test accuracy when using no augmentation. The best FID performance is achieved by no augmentation and random cropping with padding, which minimally distort the underlying data distribution $p_{\text{data}}$. Overall we find random cropping with padding provides the optimal balance between discriminative and generative performances.